OpenStudy (anonymous):

consider the equation y = 8 - 3x. for every increase of 1 in the x- variable, what is the change in the y- variable? a) increase of 3 b) increase of 8 c) decrease of 3 d) decrease of 8

OpenStudy (anonymous):

@AEYES13

OpenStudy (ranga):

Try to put x = 1 and x = 2 and see how y changes.

OpenStudy (anonymous):

The answer would be c since 8 is the constant and you are incresing the x by one every time so 8-3(1)=5 8-3(2)=2 8-3(3)=-1 is decreasing by 3 each time